OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-2,3)and is perpendicular to 2x-9y=-12

OpenStudy (a_clan):

perpendicular lines have slopes related as : m2 = -1/m1 since given slope m1 is 2/9; m1 is -9/2 Required equation is y=(-9/2)x + c put (-2,3) in this equation, 3=9 + c or c=-6 (which is the y intercept) hence final equation is y=(-9/2)x - 6